OpenStudy (anonymous):

Auditions for a play in New York were held with 120 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many men auditioned?

OpenStudy (inkyvoyd):

x+y=120 (x is number of men, y is number of women) 2x-3y=20

OpenStudy (anonymous):

that's not an option, could it be 40 maybe?

Directrix (directrix):

I cranked out 76 men and 44 women. I'll post my work.

Directrix (directrix):

@NicoleShaw --> for your perusal and comment Let the number of men be x. The number of women is 120 - x. 2x - 3(120 - x) = 20 2x - 360 + 3x = 20 5x = 360 +20 5x = 380 x = 76 --> Number of Men 120 - 76 = 44 --> Number of Women

OpenStudy (inkyvoyd):

sorry, that second quation should've been 3y-2x=20